OpenStudy (anonymous):

The slope of the tangent line to a curve at any point (x, y) on the curve is x/y. What is the equation of the curve if (3, 1) is a point on the curve?

OpenStudy (anonymous):

the slope is \(\frac{x}{y}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i think that means \[y'=\frac{x}{y}\] right?

OpenStudy (anonymous):

That's what i'm thinking

OpenStudy (anonymous):

in other words this is the unit circle

OpenStudy (anonymous):

I guess so

OpenStudy (anonymous):

ah dang i am wrong, if it was the unit circle then it would be \[y'=-\frac{x}{y}\]

OpenStudy (anonymous):

this one is "separable"

jimthompson5910 (jim_thompson5910):

y ' = x/y dy/dx = x/y y*dy = x*dx integrate both sides to get (y^2)/2 = (x^2)/2 + C

jimthompson5910 (jim_thompson5910):

hopefully you see what to do from here

OpenStudy (anonymous):

Ohh I think i see know. Give me a second

jimthompson5910 (jim_thompson5910):

are you able to determine the value of C?